OpenStudy (jazzyluvsu123):

Which of these fractions is between 1/3 and 2/5? A.) 18/60 B.) 16/45 C.) 1/2 D.) 7/30 Plz help show work FREE MEDAL :) !!!!

OpenStudy (destinymasha):

Start by finding the common denominator of 3 and 5 , closest one is 15 ,

OpenStudy (destinymasha):

what did you have to multiply to 3 to get 15, 5 so multiply the numerator 1 by 5 = 5 , and likewise for the 2 = 10.

OpenStudy (destinymasha):

now add them , 5/15 + 10/15 = 15/15 , now to find out the average , multiply by 1/2 , which is the same as dividing by 2

OpenStudy (destinymasha):

15/15 x 1/2 = 15/30

OpenStudy (destinymasha):

Notice that 15 is half of 30. So simplify it to 1/2

OpenStudy (destinymasha):

The answer is C 1/2